Two consecutive odd numbers are such that 7times the smaller subtracted from 9times the bigger gives 144.find the two numbers
Answer by Boreal(15235)

numbers are x and x+2
9(x+2)-7x=144
9x+18-7x=144
2x=126
x=63
The numbers are 63 and 65
9*65=585
7*63=441
144 is the difference